Shore Company reports the following information regarding its production cost.

Units produced 36,000 units
Direct labor $ 31 per unit
Direct materials $ 32 per unit
Variable overhead $ 10 per unit
Fixed overhead $ 102,920 in total


Compute product cost per unit under absorption costing.
Answer #1

Fixed overhead cost per unit=102920/36,000=$2.86 per unit(Approx)

product cost per unit under absorption costing=Direct labor+Direct materials+Variable overhead+Fixed overhead

=31+32+10+2.86

=75.86 per unit(Approx).